OpenStudy (anonymous):

Which of the following is not a rational expression? 5x 2x^2/11 5x/x+2 1/x^2+11x+30

ganeshie8 (ganeshie8):

last expression, is that 1/x^-1 + 11x + 30 ?

OpenStudy (anonymous):

oh, i meant 1/x^2+11x+30

ganeshie8 (ganeshie8):

ohhk to my knowledge all options are rational expressions as all can be written of the form p/q. pls wait and see what others say :D

OpenStudy (anonymous):

okay, because one of them can't be

OpenStudy (asnaseer):

a rational expression is any expression that can be written in the form p/q. however, note that the denominator cannot equal zero.

OpenStudy (asnaseer):

so, if for any of them you can think of a value for x that would make the denominator zero, then that expression is not rational

OpenStudy (anonymous):

im not sure which one

OpenStudy (asnaseer):

sorry - what I meant to say is that they ARE all rational, but some of them need to have conditions applied to them

OpenStudy (asnaseer):

to avoid the division by zero

OpenStudy (anonymous):

so what do i need to do?

OpenStudy (asnaseer):

e.g. for the 3rd one, x cannot equal -2

OpenStudy (asnaseer):

and for the last one:\[x^2+11x+30=(x+5)(x+6)\]so x cannot equal -5 or -6

OpenStudy (anonymous):

well then how would i figure out which one it is then?

OpenStudy (asnaseer):

so they are ALL rational expressions with the conditions added for the 3rd and 4th one

OpenStudy (anonymous):

which one should i got with then?

OpenStudy (asnaseer):

ALL

OpenStudy (anonymous):

i can only pick one

OpenStudy (asnaseer):

ALL are rational. Question asks which one is NOT rational - so I would say NONE of them is NOT rational

OpenStudy (anonymous):

i know, its multiple choice so i have to pick an answer

OpenStudy (asnaseer):

this might help you: http://tutorial.math.lamar.edu/Classes/Alg/RationalExpressions.aspx

OpenStudy (asnaseer):

this might be easier to understand: http://www.mathsisfun.com/algebra/rational-expression.html

OpenStudy (asnaseer):

are you SURE you have listed the 4 options correctly?

OpenStudy (anonymous):

ill repost them just in case.

OpenStudy (anonymous):

5x, 2x^2/11, 5x/x+2, 1/x^2+11x+30

OpenStudy (asnaseer):

so these are the options?\[a) 5x\]\[b)\frac{2x^2}{11}\]\[c)\frac{5x}{x+2}\]\[d)\frac{1}{x^2+11x+30}\]

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

well, I would still say that they are ALL rational expressions (with the conditions applied to (c) and (d) that I showed above). so maybe they made a mistake in the list given to you?

OpenStudy (asnaseer):

or maybe I am missing something? :)

OpenStudy (anonymous):

im not sure..